# PI-HMR: Towards Robust In-bed Temporal Human Shape Reconstruction with Contact Pressure Sensing

### v1

@@ 1 Introduction
Long-term and automatic in-bed monitoring draws increasing attentions in recent years for the growing need in heathcare, such as sleep studies [5], bedsore prevention [60], and detection of bed-exit and fall events [20]. The advancement of parameterized human representation (e.g. SMPL [36]) and human pose and shape estimation (HPS) technologies further furnish technical underpinning for the reconstruction and visualization of patient motions, facilitating caregivers to comprehend patients’ behavioral patterns in time. However, vision-based techniques, trained on in-lab or in-wild public datasets, fail in in-bed scenarios for more challenges are raised like poor illumination, occlusion by blankets, domain gaps with existing datasets (e.g. 3DPW [52]), and privacy issues in both at-home or ICUs.
Our intuition lies in that tactile serves as a crucial medium for human perception of the surroundings. Especially for in-bed scenarios, lying postures prompts full engagement between humans and environment; simultaneously, this tactile perception also encompasses valuable information about their physiques. Reconstructing human motions from these tactile feedback might provide a privacy-preserving solution to automatic in-bed management for patients and elders. Thus, many efforts have been devoted to capturing the contact pressure with a pressure-sensing bedsheet, which integrates a pressure-sensitive sensor array and collects matrix-formatted pressure distribution (named pressure images), and exploring potentials of full-body human reconstruction from these tactile sensors [8, 9, 49]. However, current methods are often constrained by model design, dataset diversity and label quality. The limitations can be categorized into three points:
@@
(3) Limited 3D label quality. One main factor limiting the data diversity is the challenge of acquiring accurate 3D labels, especially for a in-bed setting. Currently, only SLP [35] and TIP [55] datasets offer both SMPL pseudo-ground truth (p-GTs) and RGB images, with annotations in TIP being seriously doubted by depth ambiguity and penetrations due to monocular SMPLify-based optimization (in Fig. 2). Limited label quality might lead the model to misinterpret pressure cues, thus calling for a low-cost and accurate label annotation approach for in-bed scenes.
To tackle aforesaid disparities, in this work, we present a general framework bridging from annotations, model design and evaluation for pressure-based in-bed HPS tasks. Concretely, we firstly present PI-HMR, a pressure-based in-bed human shape estimation network to predict human motions from pressure sequences, as a preliminary exploration to utilize pressure characteristics. Our core philosophy falls that both joint positions and contours of high-pressure areas are essential to sense pressure distribution and its variation patterns from the redundant zero-value backgrounds. Thus, we achieve this by explicitly introducing these semantic cues, compelling the model to focus on core regions by feature sampling. Furthermore, considering that the sensing mattress is often fixed in the environment, we leverage these positional priors and feeding them into the model to learn the spatial relationship between humans and sensors. Experiments show that PI-HMR brings 17.01mm MPJPE decrease compared to PI-Mesh [55] and outperforms vision-based temporal SOTA architecture TCMR [7] (re-trained on pressure images) with 4.91mm MPJPE improvement.
@@
We highlight our key contributions: (1) a general framework for pressure-based in-bed human shape estimation task, spanning from label generation to algorithm design. (2) PI-HMR, a temporal nework to directly predict 3D meshes from in-bed pressure image sequences and outperforms both SOTA pressure-based and vision-field architectures. (3) SMPLify-IB, a gravity-based optimization technique to generate reliable SMPL p-GTs for monocular in-bed scenes. Based on SMPLify-IB, we re-generate 3D annotations for a public dataset, TIP, providing higher-quality SMPL p-GTs and mitigating implausible limb lifts due to depth ambiguity. (4) We explore the feasibility of prior expansion with knowledge distillation and TTO strategy.
@@ 2 Related Work
In-bed human pose and shape estimation. Compared with other human-related tasks, in-bed HPS faces more serious challenges from data quality and privacy issues. Thus, efforts are devoted to pursue environmental sensors for in such a non-light-of-sight (NLOS) scenes, such as infrared camera [33, 34], depth camera [16, 1, 9], pressure-sensing mattresses [8, 9, 10, 55]. Specifically for pressure-based approaches, Clever et al. [9] conducted pioneering studies by involving pressure estimation to reconstruct in-bed shapes from a single pressure image [9]. Wu et al. [55] collected a three-modality in-bed dataset TIP, and employed a VIBE-based network to predict in-bed motions from pressure sequences. Yin et al. [58] proposed a pyramid scheme to infer in-bed shapes from aligned depth, LWIR, RGB, and pressure images, and Tandon et al. [49] further improve accuracy on SLP dataset [35] with only depth and pressure modalities by integrating a pressure prediction module as auxiliary supervision.
@@ 3.2.1 Gravity Constraint
To tackle the implausible limb lifts, our rationale lies in the observations that when a person lies in bed, it should stay relaxed. Conversely, when limbs are intentionally lifted, a torque is generated at the shoulders or hips, thus resulting in discomfort. Such a conflict inspires us that when a person is motionless, all limbs should receive support to avoid a ”uncomfortable” posture. Based on such an intuition, we propose a zero-velocity detection algorithm to detect implausible limb suspensions caused by depth ambiguity and exert gravity constraints to push them into contact with the bed plane or other body parts for support. Specifically, we use velocities of 2D keypoint ground-truths to calibrate limb status. For those velocities exceeding a pre-defined threshold, we consider them to be in normal movement states; for limbs raised but nearly static, we annotate them as miscalculations from depth ambiguity and punish their distance to the bed plane. The loss term is as follows:
@@ 3.2.2 Potential-based self-penetration Constraint
Figure 3: (a) and (b): demos of our detection algorithm. $S$ is the segment, $C$ is its segment center. $v_{i}$ are vertices that need to be checked for penetration with $S$, and $v_{j}$ are the vertices from $S$ that are closest to $v_{i}$, respectively. When $\overrightarrow{v_{i}v_{j}}\cdot\overrightarrow{v_{i}C}<0$, $v_{i}$ is in penetration, and vice versa. (c) is our segments.
Figure 4: An overview of PI-HMR. We employ a multi-scale feature fusion module to enhance feature quality extracted by ResNet50
@@
To construct approximately-convex segments, we pre-define 24 segment centers (including 16 SMPL joints and 8 virtual joints in joint-sparse limbs like arms and legs to ensure uniformity), and employ a clustering algorithm to determine the assignment of SMPL vertices to segments. Finally 24 segments are generated and visualized in Fig. 3(c).
@@
$\small\mathcal{L}_{push}=|\mathbb{D}(S_{i},S_{j})|\exp(-\lambda_{pull}||c_{i}-c_{j}||)$ (2)
@@
$\small\mathcal{L}_{pull}=-|\mathbb{C}(S_{i},S_{j})|\exp(-\lambda_{push}||c_{i}-c_{j}||)$ (3)
@@ 4.1 PI-HMR
Fig. 4 shows the framework of PI-HMR. The detailed structures of each module will be elaborated as follows:
4.1.1 Overall Pipeline
Given an input pressure image sequence $V=\{I_{i}\in\mathbb{R}^{H\times W}\}^{T}_{t=1}$ with $T$ frames, PI-HMR outputs the SMPL predictions of the mid-frame by a three-stage feature extraction and fusion modules. Following [27, 7, 54], we first use ResNet50 to extract the static feature of each frame to form a static representation sequence $X=\{x_{t}\in\mathbb{R}^{2048\times H_{1}\times W_{1}}\}^{T}_{t=1}$. The extracted $X$ is then fed into our Multi-scale Feature Fusion module (MFF) to generate the fusion feature sequences $G=\{g_{t}\}^{T}_{t=1}$, with two-layer Transformer blocks behind to learn their long-term temporal dependencies and yield the temporal feature sequence $Z=\{z_{t}\}^{T}_{t=1}$. Finally, We use the mean feature of $Z$ as the integrated feature representation of the mid-frame and produce final estimations with a IEF SMPL regressor [24].
@@ 4.1.2 Multi-Scale Feature Fusion Module
where $\odot$ is the Hadamard product. The product result will be stacked with the TopK-mask and fed into a 1-layer convolution layer and Softmax layer to generate the attention matrix $H^{LK}_{i}\in\mathbb{R}^{H\times W}$. We aim to explicitly integrate these pressure distribution to enhance learnable masks’ quality. The K is set as 128 in PI-HMR, and we also conduct ablations to discuss the selection of K in Tab. 4.
@@ 4.1.3 Training Strategy
4.1.4 Encoder pre-train by cross-modal KD
@@
Method Input Modalities MPJPE PA-MPJPE MPVE Accel ACC-ERR
HMR [24] single Pressure 75.06 57.97 89.11 28.37 31.52
HMR-KD 66.30 52.41 83.01 21.23 24.41
BodyMap-WS [49] 71.48 40.91 80.08 26.62 27.98
TCMR [7] sequence 64.37 46.76 74.66 16.48 20.12
MPS-NET [54] 160.59 112.12 187.13 24.82 28.73
PI-Mesh [55] 76.47 54.65 90.54 18.44 21.86
PI-HMR (ours) 59.46 44.53 69.92 6.08 9.12
PI-HMR + KD (ours) 57.13 42.98 67.22 7.04 9.84
PI-HMR + TTO (ours) 57.76 43.31 67.76 7.01 9.83
PI-HMR + KD + TTO (ours) 55.50 41.81 65.15 7.21 9.96
@@ 4.2 Test-Time Optimization
We also explore a TTO routine to further enhance prediction quality of PI-HMR. Considering that there hasn’t been a general 2D keypoint regressor for pressure images, we are inclined towards seeking an unsupervised, prior-based optimization strategy. We notice that humans exhibit similar movement patterns across various postural states (e.g., timing, which hand to support, and leg movements). This inspires us to pre-learn such a motion habit as motion prior, playing as supplement cues to refine PI-HMR’s prediction.
We apply a VQ-VAE as the motion prior learner. The selection is rooted in our assumption that the distribution of bed-bound movements is rather constrained. In that case, for a noised motion predictions, VQ-VAE could match it to the closest pattern, thereby re-generate habit-based results. The VQ-VAE is based on Transformer blocks and show similar architecture with [13]. During training, we only auto-reconstruct the pose sequences ($\theta$ in SMPL). More details are provided in Supplementary Materials.
The VQ-VAE will act as the only motion prior and supervision in our TTO routine. For terminological convenience, given a VQ-VAE $\mathbb{M}$ and PI-HMR initial predictions $\Theta^{n}_{0}=\{\theta^{n}_{1},...\theta^{n}_{T}\}$, the $i_{th}$ iteration objectives follows:
$\small\mathcal{L}_{TTO}^{i}=\mathcal{L}_{m}(\Theta^{n}_{i},\Theta^{n}_{0})+\mathcal{L}_{m}(\Theta^{n}_{i},\mathbb{M}(\Theta^{n}_{i}))+\mathcal{L}_{sm}(\Theta^{n}_{i})$ (9)
@@ 5 Experiments
We evaluate PI-HMR on the TIP dataset. Following [55], we choose the second-to-last group of each subject as the val. set, the last group of each subject as the test set, and the remains as the training set. For evaluation, We use standard evaluation metrics including MPJPE (without pelvis alignment), PA-MPJPE, MPVE for shape errors, and Acceleration (Accel) and Acceleration errors (ACC-ERR) to evaluate smoothness. The first three metrics are measured in millimeters ($mm$), and the rest ones are measured in $mm/s^{2}$.
We compare our model with previous SOTAs and vison-based classic structures, including: HMR [24] and HMR-KD (HMR structure with and without cross-modal KD), BodyMap-WS [49], TCMR [7], MPS-NET [54], and PI-Mesh [55]. All methods are re-trained on TIP with our re-generated SMPL p-GTs, and follows the same training setups with PI-HMR. We provide detailed implementation details of these approaches and PI-HMR in Sup. Mat.
@@ 5.1 Overall Results for PI-HMR
We present quantitative evaluations in Tab. 1. Our methods outperform all image or sequence-based methods, presenting about 17.01mm MPJPE decrease compared to PI-Mesh and also outperforms SOTA vision-based architecture HMR, TCMR with 15.6mm, 4.91mm MPJPE improvement, while maintaining comparable ACC-ERR compared with SOTA approaches. Moreover, our introduced cross-modal KD and TTO strategy further improve the robustness of PIHMR, brings 2.33mm and 1.7mm MPJPE improvements compared with basic structure. In particular, the TTO strategy, as an unsupervised, entirely prior-based optimization strategy, demonstrates the effectiveness of learning and refinement based on user habits. We provide visual comparisons between CLiff, PI-Mesh and PI-HMR in Fig. 6.

### v2

@@ 1 Introduction
Long-term and automatic in-bed monitoring draws increasing attention in recent years for the growing need in heathcare, such as sleep studies [5], bedsore prevention [60], and detection of bed-exit and fall events [20]. The advancement of parameterized human representation (e.g. SMPL [36]) and human pose and shape estimation (HPS) technologies further furnish technical underpinning for the reconstruction and visualization of patient motions, facilitating caregivers to comprehend patients’ behavioral patterns in time. However, vision-based techniques, trained on in-lab or in-wild public datasets, fail in in-bed scenarios for more challenges are raised like poor illumination, occlusion by blankets, domain gaps with existing datasets (e.g. 3DPW [52]), and privacy issues in both at-home or ICUs.
Our intuition lies in that tactile serves as a crucial medium for human perception of the surroundings. Especially for in-bed scenarios, lying postures prompt full engagement between humans and environment; simultaneously, this tactile perception also encompasses valuable information about their physiques. Reconstructing human motions from this tactile feedback might provide a privacy-preserving solution to automatic in-bed management for patients and elders. Thus, many efforts have been devoted to capturing the contact pressure with a pressure-sensing bedsheet, which integrates a pressure-sensitive sensor array and collects matrix-formatted pressure distribution (named pressure images), and exploring potentials of full-body human reconstruction from these tactile sensors [8, 9, 49]. However, current methods are often constrained by model design, dataset diversity and label quality. The limitations can be categorized into three points:
@@
(3) Limited 3D label quality. One main factor limiting the data diversity is the challenge of acquiring accurate 3D labels, especially for an in-bed setting. Currently, only SLP [35] and TIP [55] datasets offer both SMPL pseudo-ground truth (p-GTs) and RGB images, with annotations in TIP being seriously doubted by depth ambiguity and penetrations due to monocular SMPLify-based optimization (in Fig. 2). Limited label quality might lead the model to misinterpret pressure cues, thus calling for a low-cost and accurate label annotation approach for in-bed scenes.
To tackle aforesaid disparities, in this work, we present a general framework bridging from annotations, model design and evaluation for pressure-based in-bed HPS tasks. Concretely, we firstly present PI-HMR, a pressure-based in-bed human shape estimation network to predict human motions from pressure sequences, as a preliminary exploration to utilize pressure characteristics. Our core philosophy falls that both joint positions and contours of high-pressure areas are essential to sense pressure distribution and its variation patterns from the redundant zero-value backgrounds. Thus, we achieve this by explicitly introducing these semantic cues, compelling the model to focus on core regions by feature sampling. Furthermore, considering that the sensing mattress is often fixed in the environment, we leverage these positional priors and feed them into the model to learn the spatial relationship between humans and sensors. Experiments show that PI-HMR brings 17.01mm MPJPE decrease compared to PI-Mesh [55] and outperforms vision-based temporal SOTA architecture TCMR [7] (re-trained on pressure images) with 4.91mm MPJPE improvement.
@@
We highlight our key contributions: (1) a general framework for pressure-based in-bed human shape estimation task, spanning from label generation to algorithm design. (2) PI-HMR, a temporal network to directly predict 3D meshes from in-bed pressure image sequences and outperforms both SOTA pressure-based and vision-field architectures. (3) SMPLify-IB, a gravity-based optimization technique to generate reliable SMPL p-GTs for monocular in-bed scenes. Based on SMPLify-IB, we re-generate 3D annotations for a public dataset, TIP, providing higher-quality SMPL p-GTs and mitigating implausible limb lifts due to depth ambiguity. (4) We explore the feasibility of prior expansion with knowledge distillation and TTO strategy.
@@ 2 Related Work
In-bed human pose and shape estimation. Compared with other human-related tasks, in-bed HPS faces more serious challenges from data quality and privacy issues. Thus, efforts are devoted to pursuing environmental sensors for in such a non-light-of-sight (NLOS) scenes, such as infrared camera [33, 34], depth camera [16, 1, 9], pressure-sensing mattresses [8, 9, 10, 55]. Specifically for pressure-based approaches, Clever et al. [9] conducted pioneering studies by involving pressure estimation to reconstruct in-bed shapes from a single pressure image [9]. Wu et al. [55] collected a three-modality in-bed dataset TIP, and employed a VIBE-based network to predict in-bed motions from pressure sequences. Yin et al. [58] proposed a pyramid scheme to infer in-bed shapes from aligned depth, LWIR, RGB, and pressure images, and Tandon et al. [49] improves accuracy on SLP [35] with depth and pressure modalities by integrating a pressure prediction module as auxiliary supervision.
@@ 3.2.1 Gravity Constraint
To tackle the implausible limb lifts, our rationale lies in the observations that when a person lies in bed, it should stay relaxed. Conversely, when limbs are intentionally lifted, a torque is generated at the shoulders or hips, thus resulting in discomfort. Such a conflict inspires us that when a person is motionless, all limbs should receive support to avoid an "uncomfortable" posture. Based on such an intuition, we propose a zero-velocity detection algorithm to detect implausible limb suspensions caused by depth ambiguity and exert gravity constraints to push them into contact with the bed plane or other body parts for support. Specifically, we use velocities of 2D keypoint ground-truths to calibrate limb status. For those velocities exceeding a pre-defined threshold, we consider them to be in normal movement states; for limbs raised but nearly static, we annotate them as miscalculations from depth ambiguity and punish their distance to the bed plane. The loss term is as follows:
@@ 3.2.2 Potential-based self-penetration Constraint
Figure 3: (a) and (b): demos of our detection algorithm. $S$ is the segment, $C$ is its segment center. $v_{i}$ are vertices that need to be checked for penetration with $S$, and $v_{j}$ are the vertices from $S$ that are closest to $v_{i}$, respectively. When $\overrightarrow{v_{i}v_{j}}\cdot\overrightarrow{v_{i}C}<0$, $v_{i}$ is in penetration, and vice versa. (c) is our segment.
Figure 4: An overview of PI-HMR. PI-HMR outputs the midframe’s SMPL predictions of the whole sequence.
@@
To construct approximately-convex segments, we pre-define 24 segment centers (including 16 SMPL joints and 8 virtual joints in joint-sparse limbs like arms and legs to ensure uniformity), and employ a clustering algorithm to determine the assignment of SMPL vertices to segments. Finally, 24 segments are generated and visualized in Fig. 3(c).
@@
$\small\mathcal{L}_{push}=|\mathbb{D}(S_{i},S_{j})|\exp(-\lambda_{push}||c_{i}-c_{j}||)$ (2)
@@
$\small\mathcal{L}_{pull}=-|\mathbb{C}(S_{i},S_{j})|\exp(-\lambda_{pull}||c_{i}-c_{j}||)$ (3)
@@ 4.1 PI-HMR
Our motivation is to utilize pressure data nature. So our efforts fall into three stages: alleviating the dataset bottleneck and learning cross-dataset human and motion priors in the pre-training stage; pressure-based PI-HMR’s design; and learning user’s habits to overcome information ambiguity in the TTO. Thus, the data flow includes: (1) pre-train: KD-based pre-training with the training set; (2) train: train the PI-HMR and VQ-VAE with the training set; (3) test: test with PI-HMR on the test set and improve the estimates with the TTO strategy. Fig. 4 shows the framework of PI-HMR. The details of each module will be elaborated as follows:
4.1.1 Overall Pipeline of PI-HMR
Given an input pressure image sequence $V=\{I_{i}\in\mathbb{R}^{H\times W}\}^{T}_{t=1}$ with $T$ frames, PI-HMR outputs the SMPL predictions of the mid-frame by a three-stage feature extraction and fusion modules. Following [27, 7, 54], we first use ResNet50 to extract the static feature of each frame to form a static representation sequence $X=\{x_{t}\in\mathbb{R}^{2048\times H_{1}\times W_{1}}\}^{T}_{t=1}$. The extracted $X$ is then fed into our Multi-scale Feature Fusion module (MFF) to generate the fusion feature sequences $G=\{g_{t}\}^{T}_{t=1}$, with two-layer Transformer blocks behind to learn their long-term temporal dependencies and yield the temporal feature sequence $Z=\{z_{t}\}^{T}_{t=1}$. Finally, We use the mean feature of $Z$ as the integrated feature representation of the mid-frame and produce final estimations with an IEF SMPL regressor [24].
@@ 4.1.2 Multi-Scale Feature Fusion Module
where $\odot$ is the Hadamard product. The product result will be stacked with the TopK-mask and fed into a 1-layer convolution layer and Softmax layer to generate the attention matrix $H^{LK}_{i}\in\mathbb{R}^{H\times W}$. We aim to explicitly integrate these pressure distributions to enhance learnable masks’ quality. The K is set as 128 in PI-HMR, and we also conduct ablations to discuss the selection of K in Tab. 4.
Auxiliary Joint Regressor. We use an auxiliary joint regressor to provide 2D joints for the multi-scale feature extraction (shown in Fig. 5). The regressor takes the static feature $x_{i}$ as input and returns the 2D positions of 12 joints in the pressure image, noted as $J_{i}^{2D}$. The 2D regressor will be trained in conjunction with the entire model.
@@ 4.1.3 Training Strategy
4.2 Encoder pre-train by cross-modal KD
@@
Method Input Modalities MPJPE PA-MPJPE MPVE ACC-ERR
HMR [24] single Pressure 75.06 57.97 89.11 31.52
HMR-KD 66.30 52.41 83.01 24.41
BodyMap-WS [49] 71.48 40.91 80.08 27.98
TCMR [7] sequence 64.37 46.76 74.66 20.12
MPS-NET [54] 160.59 112.12 187.13 28.73
PI-Mesh [55] 76.47 54.65 90.54 21.86
PI-HMR (ours) 59.46 44.53 69.92 9.12
PI-HMR + KD (ours) 57.13 42.98 67.22 9.84
PI-HMR + TTO (ours) 57.76 43.31 67.76 9.83
PI-HMR + KD + TTO (ours) 55.50 41.81 65.15 9.96
@@ 4.3 Test-Time Optimization
We also explore a TTO routine to further enhance prediction quality of PI-HMR. Considering that there hasn’t been a general 2D keypoint regressor for pressure images, we are inclined toward seeking an unsupervised, prior-based optimization strategy. We notice that humans exhibit similar movement patterns across various postural states (e.g., timing, which hand to support, and leg movements). This inspires us to pre-learn such a motion habit as motion prior, playing as supplement cues to refine PI-HMR’s prediction.
We apply a VQ-VAE as the motion prior learner. The selection is rooted in our assumption that the distribution of bed-bound movements is rather constrained. In that case, for a noised motion prediction, VQ-VAE could match it to the closest pattern, thereby re-generating habit-based results. The VQ-VAE is based on Transformer blocks and show similar architecture with [13]. During training, we only auto-reconstruct the pose sequences ($\theta$ in SMPL). More details are provided in Supplementary Materials.
The VQ-VAE will act as the only motion prior and supervision in our TTO routine. For terminological convenience, given a VQ-VAE $\mathbb{M}$ and PI-HMR initial predictions $\Theta^{0}=\{\theta^{0}_{1},...\theta^{0}_{T}\}$, the $i_{th}$ iteration objectives follows:
$\small\mathcal{L}_{TTO}^{i}=\mathcal{L}_{m}(\Theta^{i},\Theta^{0})+\mathcal{L}_{m}(\Theta^{i},\mathbb{M}(\Theta^{i}))+\mathcal{L}_{sm}(\Theta^{i})$ (9)
@@ 5 Experiments
We evaluate PI-HMR on the TIP dataset. Following [55], we choose the second-to-last group of each subject as the val. set, the last group of each subject as the test set, and the remains as the training set. For evaluation, We use standard evaluation metrics including MPJPE (without pelvis alignment), PA-MPJPE, MPVE for shape errors, and Acceleration errors (ACC-ERR) to evaluate smoothness. The first three metrics are measured in millimeters ($mm$), and the rest are measured in $mm/s^{2}$.
We compare our model with previous SOTAs and vison-based classic structures, including: HMR [24] and HMR-KD (HMR structure with and without cross-modal KD), BodyMap-WS [49], TCMR [7], MPS-NET [54], and PI-Mesh [55]. All methods are re-trained on TIP with our re-generated SMPL p-GTs, and follow the same training setups with PI-HMR. We provide detailed implementation details of these approaches and PI-HMR in Sup. Mat.
@@ 5.1 Overall Results for PI-HMR
We present quantitative evaluations in Tab. 1. Our methods outperform all image or sequence-based methods, presenting about 17.01mm MPJPE decrease compared to PI-Mesh and also outperforms SOTA vision-based architecture HMR, TCMR with 15.6mm, 4.91mm MPJPE improvement, while maintaining comparable ACC-ERR compared with SOTA approaches. Moreover, our introduced cross-modal KD and TTO strategy further improve the robustness of PIHMR, bringing 2.33mm and 1.7mm MPJPE improvements compared with basic structure. In particular, the TTO strategy, as an unsupervised, entirely prior-based optimization strategy, demonstrates the effectiveness of learning and refinement based on user habits. We provide visual comparisons between CLIFF, PI-Mesh and PI-HMR in Fig. 6.